Two cars leave from Ohio and go in opposite directions. One car goes west at 60 mph. Another goes east at 80 mph but leaves 1 hour 30 min after the first car. At what time will they be 620 miles apart? This isn't from a book. I am completely stumped with no idea how to do this.
Answer by checkley77(12569)

LET 60*T= THE WEST CAR.
LET 80(T-1.5)= THE EAST CAR.
60T+80(T-1.5)=620
60T+80T-120=620
140T=620+120
140T=740
T=740/140
T=5.2857 HOURS AFTER THE CAR GOING WEST STARTED THEY WILL BE 620 MILES APART
PROOF:
60*5.2857+80(5.2857-1.5)=620
317.14285+80*3.7857=620
317.14285+302.856=620
620~620